# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v2

@@ 6.2  Computing a conforming sliding sequence $\sigma$
Every time we increment $i$ or $j$, we update $\mathcal{E}[i-1,j+1]$ accordingly in $O(\log n)$ time. To decide whether to increment $i$ or $j$, we invoke Lemma˜17 using $s=i$ and $t=j$ to check if there exists a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]$. If yes, then we add the pair $(i,j)$ to $\sigma^{*}$ and increment $j$. Otherwise, we add the pair (i+1,j1)(i+1,j-1) to $\sigma^{*}$ and increment $i$.
Correctness.
If we start our sliding window at $(i,j)=(1,2)$ then our invariants immediately imply that when we reach $i=n+1$, the sequence $\sigma^{*}$ have been computed correctly in $O(n\log n)$ time. What remains is to prove that the invariants always hold. Clearly, the first invariant is always maintained. The second invariant is maintained by construction because before we increment $j$, we add $(i,j)$ to $\sigma^{*}$ and before we increment $i$, we add (i+1,j1)(i+1,j-1) to $\sigma^{*}$.
Next, consider the third invariant. If we are about to increment $j$, then exists a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]\subset\mathcal{E}[i-1,j]$, thus after incrementing $j$ the third invariant still holds. Suppose otherwise that we are about to increment $i$. By the third invariant, there exists a guard (g,[i,j1])(g,[i,j-1]) with $g\in\mathcal{E}[i-1,j]$. This guard also guards $[i+1,j]$ and $g\in\mathcal{E}[i-1,j]\subset\mathcal{E}[i,j]$. So, the guard (g,[i,j1])(g,[i,j-1]) implies that the third invariant is maintained after incrementing $i$.
Finally, we consider the fourth invariant. Before we are about to increment $j$, $(i,j)$ is added to $\sigma^{*}$. Thus, if the fourth invariant holds for up to $(i,j)$ then it must hold after adding $(i,j)$ to $\sigma^{*}$ and incrementing $j$. If we are about to increment $i$ then there does not exist a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]$. Suppose for the sake of contradiction that there exists a guard $(g,[i,j^{\prime}])$ with $g\in\mathcal{E}[i-1,j^{\prime}+1]$ and $j^{\prime}>j$. However, $g\in\mathcal{E}[i-1,j^{\prime}+1]\subset\mathcal{E}[i-1,j+1]$ and $i,j]\subset[i,j^{\prime}]$ – a contradiction. We now apply the third invariant to note that there exists a guard (g,[i,j1])(g,[i,j-1]) with $g\in\mathcal{E}[i-1,j]$. The fact that there exists no guard $(g,[i,j^{\prime}])$ with $g\in\mathcal{E}[i-1,j^{\prime}+1]$ and $j^{\prime}>j$, and one guard (g,[i,j1])(g,[i,j-1]) with $g\in\mathcal{E}[i-1,j]$, implies that implies that $d_{\text{max}}(i)=j-1$. It follow that there is no pair $(i,j^{\prime})$ with $j^{\prime}\geq j$ for which $j^{\prime}\in[i,d_{\max}(i^{\prime})]$, and we can safely increment $i$. ∎
@@ 6.3  A linear-size set $\mathcal{B}_{\sigma}$ that contains all reduced good dominators in $\mathcal{B}$
We compute $\mathcal{B}_{\sigma}$ by computing for all $x\in V_{\sigma}$ some maximal chain $\ell_{x},r_{x}]$, and applying Lemma˜13.
Lemma 19.
@@
Proof.
We iterate over all $(i,j)\in\sigma$ in order. This requires $O(n)$ updates, where each update increments either $i$ or $j$. Hence, we can maintain the vertices of $\mathcal{E}[i-1,j+1]$ explicitly in memory using the first-in-first-out data structure from Corollary˜11. By Lemma˜10, each point appears as a vertex of $\mathcal{E}[i-1,j+1]$ exactly once, and once removed, it never reappears. During this update sequence, we maintain a set of active points: points in the plane that are not necessarily current vertices of $\mathcal{E}[i-1,j+1]$ but are still under consideration. Whenever we advance in $\sigma$, we add all new vertices of $\mathcal{E}[i-1,j+1]$ to the active set. By construction, no point is ever added twice.
@@
By Lemma˜9 and Lemma˜10, at most $O(n)$ points are ever added to the active set, and each point is removed at most once. As each insertion, deletion, and query takes $O(\log n)$ time, the total running time is $O(n\log n)$. ∎
Corollary 20.
@@
Proof.
This is an immediate consequence of Lemma˜19 as we can iterate over $\sigma$ once forward, and once backward. ∎
Theorem 21.
@@
Proof.
Lemma˜9 implies that $|V_{\sigma}|\in O(n)$, so $|\mathcal{B}_{\sigma}|\in O(n)$. We apply Corollary˜20 and compute for each $x\in V_{\sigma}$ the chain $\ell_{x},r_{x}]$. We then apply for each $x\in V_{\sigma}$ Lemma˜13, which yields $\mathcal{B}_{\sigma}$ in $O(n\log n)$ time. We next prove that $\mathcal{B}_{\sigma}$ contains the set of reduced bad dominators. Suppose for contradiction that there is a reduced bad dominator $(c,[u_{\text{max}},v_{\text{max}}])$ that is not in $\mathcal{B}_{\sigma}$. Then $c$ is a vertex of some $\mathcal{E}[i-1,j+1]$, defined by two edges [a1,a][a-1,a] and $[b,b+1]$, and $u_{\max}\in[i-1,i)$ and $v_{\max}\in(j,j+1]$. If $(i,j)\in\sigma$, then $(c,[u_{\text{max}},v_{\text{max}}])$ would be a candidate guard and thus in $\mathcal{B}_{\sigma}$. So, $(i,j)\notin\sigma$. By definition of $\sigma$, it must be that $\textnormal{{next}}(u_{\text{max}})>j+1$. However, by definition of the reduced bad dominators (Definition˜6) there is no guard dominating $(c,[u_{\text{max}},v_{\text{max}}])$, contradicting that $\textnormal{{next}}(u_{\text{max}})>j+1$. We conclude all reduced dominators are in $\mathcal{B}_{\sigma}$.
To construct the corresponding querying data structure, we loop over all $(i,j)\in\sigma$ in order, which discretizes time $t\in[0,|\sigma|]$ where at time $t$ our loop is at $(i_{t},j_{t})\in\sigma$. For each $g\in V_{\sigma}$ there is a unique time interval $t_{1},t_{2}]$ where for all $t\in[t_{1},t_{2}]$, $g$ is a vertex of $\mathcal{E}[i_{t}-1,j_{t}+1]$. We loop over all $t$ and use the first-in-first-out convex hull data structure from Corollary˜11 to maintain $\mathcal{E}[i_{t}-1,j_{t}+1]$ explicitly in $O(n\log n)$ time and $O(n)$ space. This computes for all $g\in V_{\sigma}$ their time interval $t_{1},t_{2}]$. Let $g\in V_{\sigma}$, $(g,[u_{\text{max}},v_{\text{max}}])\in\mathcal{B}_{\sigma}$ and $t_{1},t_{2}]$ be the corresponding interval. We create a weighted rectangle $R_{g}:=[t_{1},t_{2}]\times[u_{\text{max}},v_{\text{max}}]$ in $\mathbb{R}^{2}$ where the weight is $v_{\text{max}}$. This creates a set $R$ of $O(n)$ weighted rectangles in the plane, which we store in a stabbing-query data structure that for any query point $q$, returns the maximum-weighted rectangle in $R$ that intersects $q$ in logarithmic time. Such a stabbing query data structure of size $O(n)$ can be implemented in various ways through standard techniques (for details, we refer to the stabbing query implementations in [Agarwal2005Stabbing]). Given a query point $x\in\partial P$ with $(i,j)\in\sigma$, we compute in $O(\log n)$ the corresponding time $t$ such that $(i,j)=(i_{t},j_{t})$. We then perform a stabbing query with the point $(t,x)$ and the maximum-weight rectangle corresponds to the desired query output. ∎
@@ 7  An $O(kn\log n)$-time algorithm
We will now present an algorithm, which for a set $X\subset\partial P$ will compute $\textnormal{{next}}(x)$ for every $x\in X$ in total time $O((n+|X|)\log n)$. This then yields a $O(kn\log n)$ algorithm by applying this subroutine $k$ times, together with Theorem˜7. Figures 14 and 14 show how the next function is recursively applied to different starting points in the same polygon.
Lemma 22.
Let $P$ be a polygon with $n$ vertices. Let $X\subset\partial P$ be given with $|X|\in O(n)$. There is an algorithm which computes a guard $(g,[x,\textnormal{{next}}(x)])$ for every $x\in X$ in total time $O((n+|X|)\log n)$ using $O(n)$ space.
Proof.
As a preprocessing step, we compute in $O(n\log n)$ time via Theorems 16, 18 and 21:
@@
  • a segment tree $T$ that stores, for all $(g,[u,v])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$, the interval $[u,v]$ (together with $g$).
@@
Let $x$ denote the current head of the queue. For any pair $(i^{\prime},j^{\prime})\in\sigma$ for which $x\in[i^{\prime},i^{\prime}+1)$, there exists a guard $(g,[x,j^{\prime}])$ by definition of $\sigma$. If $x\notin[i-1,i)$, we continue traversing $\sigma$ (incrementing either $i$ or $j$). We are guaranteed that during this traversal, $j\leq\textnormal{{next}}(x)$. Once $(i,j)$ is such that $x\in[i-1,i)$, we have $\textnormal{{next}}(x)\in(j,j+1]$ if and only if there exists no contiguous guard $(g,[x,j+1])$ with $g\in\mathcal{E}[i-1,j+2]$. We therefore apply Lemma˜17 using the lookahead queries of the Sliding-Window data structure (3) to test in $O(\log n)$ time whether a guard $(g,[x,j+1])$ with $g\in\mathcal{E}[i-1,j+2]$ exists, and, if so, we increment $j$. This way, we find the pair $(i,j)$ that corresponds to $x,\textnormal{{next}}(x)$.
@@
We compute the shortest path from $x$ to $j+1$ in $O(\log n)$ time and obtain the supporting line $\ell$ of its first edge. A ray-shooting query in Sliding-Window data structure (3) identifies the last point of intersection $g^{*}$ between $\ell$ and $\mathcal{E}[i-1,j+1]$. By definition, the only ugly dominator that sees $x,\textnormal{{next}}(x)]$ has $g^{*}$ as its guard. We apply Lemma˜13 to compute the maximal chain $u_{\max},v_{\max}]$ that contains $j$ and is visible from $g^{*}$ in $O(\log n)$ time.
It follows that $\textnormal{{next}}(x)$ is realized either by $(g,[u,v])$ or by $(g^{*},[u_{\max},v_{\max}])$, and we can distinguish between these cases in constant time. We dequeue $x$ and proceed to the next element in the queue. Each $O(\log n)$ step is charged either to an increment of $j$ or to the removal of one element from the queue. Hence, the total running time is $O((n+|X|)\log n)$. ∎
Theorem 23.
@@
Proof.
Let $X$ be the set that contains all vertices of $P$ and includes for all guards $(g,[u,v])\in\mathcal{D}^{\prime}$ the point $u$. Theorem˜16 implies $X$ can be computed in $O(n\log n)$ time. By Theorem˜7 and the proof of Theorem˜8, it follows that if we apply $\textnormal{{next}}(x)$ to each $x\in X$, recursively, $k$ times then for at least one $x\in X$ the result is a point $v>x+n$. In other words, for this $x\in X$ the recursive application of $\textnormal{{next}}(x)$ results in a solution of size $k$ that guards $\partial P$.
We first compute $k$ and such a corresponding $x\in X$, and afterwards find the corresponding optimal guarding solution. Via Lemma˜22, we can compute $\textnormal{{next}}(x)$ for every $x$ in $X$. We do this repeatedly, computing, and maintaining, $X_{i}=\{\textnormal{{next}}(x)|x\in X_{i-1}\}$, with $X_{0}=X$, until there is a point $x$ such that $\textnormal{{next}}^{k}(x)\geq x+n$. This takes $k$ rounds, and thus a total of $O(kn\log n)$ time, and returns the size $k$ of the optimal solution and some $x$. To also output the corresponding set of guards, we apply Lemma˜22 once more with an input $X=\{x\}$ where each time we find a guard $(g,[x,\textnormal{{next}}(x)])$, we quickly add $\textnormal{{next}}(x)$ to $X$ and the queue before we dequeue $x$. Since Lemma˜22 computes the explicit guard $(g,[x,\textnormal{{next}}(x)])$, the result are $k$ guards that together see all of $\partial P$. ∎
For the remainder of this paper we will assume that we know that there is no solution of size $3$ or less. This allows us to impose stronger structure, which in turn enables us to give a $O(n\log n)$ algorithm for a polygon where only solutions of size at least $4$ exist. To this end, we show that Theorem˜23 implies an $O(n\log n)$-time test for this property:
Corollary 24.
@@
Proof.
We simply stop the algorithm described in Theorem˜23 after $K$ rounds. ∎

### v4

@@ 6.2 Computing a conforming sliding sequence $\sigma$
Every time we increment $i$ or $j$, we update $\mathcal{E}[i-1,j+1]$ accordingly in $O(\log n)$ time. To decide whether to increment $i$ or $j$, we invoke Lemma˜6.9 using $s=i$ and $t=j$ to check if there exists a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]$. If yes, then we add the pair $(i,j)$ to $\sigma^{*}$ and increment $j$. Otherwise, we add the pair (i+1,j1)(i+1,j-1) to $\sigma^{*}$ and increment $i$.
Correctness.  If we start our sliding window at $(i,j)=(1,2)$ then our invariants immediately imply that when we reach $i=n+1$, the sequence $\sigma^{*}$ have been computed correctly in $O(n\log n)$ time. What remains is to prove that the invariants always hold. Clearly, the first invariant is maintained. The second invariant is maintained by construction because before we increment $j$, we add $(i,j)$ to $\sigma^{*}$ and before we increment $i$, we add (i+1,j1)(i+1,j-1) to $\sigma^{*}$.
Next, consider the third invariant. If we are about to increment $j$, then exists a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]\subseteq\mathcal{E}[i-1,j]$, thus after incrementing $j$ the third invariant still holds. Suppose otherwise that we are about to increment $i$. By the third invariant, there exists a guard (g,[i,j1])(g,[i,j-1]) with $g\in\mathcal{E}[i-1,j]$. This guard also guards $[i+1,j]$ and $g\in\mathcal{E}[i-1,j]\subseteq\mathcal{E}[i,j]$. So, the guard (g,[i,j1])(g,[i,j-1]) implies that the third invariant is maintained after incrementing $i$.
Finally, we consider the fourth invariant. Before we are about to increment $j$, $(i,j)$ is added to $\sigma^{*}$. Thus, if the fourth invariant holds for up to $(i,j)$ then it must hold after adding $(i,j)$ to $\sigma^{*}$ and incrementing $j$. If we are about to increment $i$ then there does not exist a guard $(g,[i,j])$ with $g\in\mathcal{E}[i-1,j+1]$. Suppose for the sake of contradiction that there exists a guard $(g,[i,j^{\prime}])$ with $g\in\mathcal{E}[i-1,j^{\prime}+1]$ and $j^{\prime}>j$. However, $g\in\mathcal{E}[i-1,j^{\prime}+1]\subseteq\mathcal{E}[i-1,j+1]$ and $i,j]\subseteq[i,j^{\prime}]$ – a contradiction. We now apply the third invariant to note that there exists a guard (g,[i,j1])(g,[i,j-1]) with $g\in\mathcal{E}[i-1,j]$. The fact that there exists no guard $(g,[i,j^{\prime}])$ with $g\in\mathcal{E}[i-1,j^{\prime}+1]$ and $j^{\prime}>j$, and one guard (g,[i,j1])(g,[i,j-1]) with $g\in\mathcal{E}[i-1,j]$, implies that implies that $d_{\text{max}}(i)=j-1$. It follow that there is no pair $(i,j^{\prime})$ with $j^{\prime}\geq j$ for which $j^{\prime}\in[i,d_{\max}(i^{\prime})]$, and we can safely increment $i$.
@@ 6.3 A linear-size set $\mathcal{B}_{\sigma}$ that contains all reduced good dominators in $\mathcal{B}$
We compute $\mathcal{B}_{\sigma}$ by computing for all $x\in V_{\sigma}$ some maximal chain $\ell_{x},r_{x}]$, and applying Lemma˜6.1.
Lemma 6.15.
@@
Proof 6.16.
We iterate over all $(i,j)\in\sigma$ in order. This requires $O(n)$ updates, where each update increments either $i$ or $j$. Hence, we can maintain the vertices of $\mathcal{E}[i-1,j+1]$ explicitly in memory using the first-in-first-out data structure from Corollary˜5.5. By Lemma˜5.3, each point appears as a vertex of $\mathcal{E}[i-1,j+1]$ exactly once, and once removed, it never reappears. During this update sequence, we maintain a set of active points: points in the plane that are not necessarily current vertices of $\mathcal{E}[i-1,j+1]$ but are still under consideration. Whenever we advance in $\sigma$, we add all new vertices of $\mathcal{E}[i-1,j+1]$ to the active set. By construction, no point is ever added twice.
@@
By Lemma˜5.2 and Lemma˜5.3, at most $O(n)$ points are ever added to the active set, and each point is removed at most once. As each insertion, deletion, and query takes $O(\log n)$ time, the total running time is $O(n\log n)$.
Corollary 6.17.
@@
Proof 6.18.
This is an immediate consequence of Lemma˜6.15 as we can iterate over $\sigma$ once forward, and once backward.
Theorem 6.19.
@@
Proof 6.20.
Lemma˜5.2 implies that $|V_{\sigma}|\in O(n)$, so $|\mathcal{B}_{\sigma}|\in O(n)$. We apply Corollary˜6.17 and compute for each $x\in V_{\sigma}$ the chain $\ell_{x},r_{x}]$. We then apply for each $x\in V_{\sigma}$ Lemma˜6.1, which yields $\mathcal{B}_{\sigma}$ in $O(n\log n)$ time. We next prove that $\mathcal{B}_{\sigma}$ contains the set of reduced bad dominators. Suppose for contradiction that there is a reduced bad dominator $(c,[u_{\text{max}},v_{\text{max}}])$ that is not in $\mathcal{B}_{\sigma}$. Then $c$ is a vertex of some $\mathcal{E}[i-1,j+1]$, defined by two edges [a1,a][a-1,a] and $[b,b+1]$, and $u_{\max}\in[i-1,i)$ and $v_{\max}\in(j,j+1]$. If $(i,j)\in\sigma$, then $(c,[u_{\text{max}},v_{\text{max}}])$ would be a candidate guard and thus in $\mathcal{B}_{\sigma}$. So, $(i,j)\notin\sigma$. By definition of $\sigma$, it must be that $\textnormal{{next}}(u_{\text{max}})>j+1$. However, by definition of the reduced bad dominators (Definition˜4.12) there is no guard dominating $(c,[u_{\text{max}},v_{\text{max}}])$, contradicting that $\textnormal{{next}}(u_{\text{max}})>j+1$. We conclude all reduced dominators are in $\mathcal{B}_{\sigma}$.
To construct the corresponding querying data structure, we loop over all $(i,j)\in\sigma$ in order, which discretizes time $t\in[0,|\sigma|]$ where at time $t$ our loop is at $(i_{t},j_{t})\in\sigma$. For each $g\in V_{\sigma}$ there is a unique time interval $t_{1},t_{2}]$ where for all $t\in[t_{1},t_{2}]$, $g$ is a vertex of $\mathcal{E}[i_{t}-1,j_{t}+1]$. We loop over all $t$ and use the first-in-first-out convex hull data structure from Corollary˜5.5 to maintain $\mathcal{E}[i_{t}-1,j_{t}+1]$ explicitly in $O(n\log n)$ time and $O(n)$ space. This computes for all $g\in V_{\sigma}$ their time interval $t_{1},t_{2}]$. Let $g\in V_{\sigma}$, $(g,[u_{\text{max}},v_{\text{max}}])\in\mathcal{B}_{\sigma}$ and $t_{1},t_{2}]$ be the corresponding interval. We create a weighted rectangle $R_{g}:=[t_{1},t_{2}]\times[u_{\text{max}},v_{\text{max}}]$ in $\mathbb{R}^{2}$ where the weight is $v_{\text{max}}$. This creates a set $R$ of $O(n)$ weighted rectangles in the plane, which we store in a stabbing-query data structure that for any query point $q$, returns the maximum-weighted rectangle in $R$ that intersects $q$ in logarithmic time. Such a stabbing query data structure of size $O(n)$ can be implemented in various ways through standard techniques (for details, we refer to the stabbing query implementations in [Agarwal2005Stabbing]). Given a query point $x\in\partial P$ with $(i,j)\in\sigma$, we compute in $O(\log n)$ the corresponding time $t$ such that $(i,j)=(i_{t},j_{t})$. We then perform a stabbing query with the point $(t,x)$ and the maximum-weight rectangle corresponds to the desired query output.
@@ 7 An $O(kn\log n)$-time algorithm
We will now present an algorithm, which for a set $X\subseteq\partial P$ will compute $\textnormal{{next}}(x)$ for every $x\in X$ in total time $O((n+|X|)\log n)$. This then yields an $O(kn\log n)$ algorithm by applying this subroutine $k$ times, together with Theorem˜4.17. Figures 14 and 14 show how the next function is recursively applied to different starting points in the same polygon.
Lemma 7.1.
Let $P$ be a polygon with $n$ vertices. Let $X\subseteq\partial P$ be given with $|X|\in O(n)$. There is an algorithm which computes a guard $(g,[x,\textnormal{{next}}(x)])$ for every $x\in X$ in total time $O((n+|X|)\log n)$ using $O(n)$ space.
Proof 7.2.
As a preprocessing step, we compute in $O(n\log n)$ time via Theorems 6.7, 6.11 and 6.19:
@@
  • a segment tree $T$ that stores, for all $(g,[u,v])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$, the interval $[u,v]$ (with $g$).
@@
Let $x$ denote the current head of the queue. For any pair $(i^{\prime},j^{\prime})\in\sigma$ for which $x\in[i^{\prime},i^{\prime}+1)$, there exists a guard $(g,[x,j^{\prime}])$ by definition of $\sigma$. If $x\notin[i-1,i)$, we continue traversing $\sigma$ (incrementing either $i$ or $j$). We are guaranteed that during this traversal, $j\leq\textnormal{{next}}(x)$. Once $(i,j)$ is such that $x\in[i-1,i)$, we have $\textnormal{{next}}(x)\in(j,j+1]$ if and only if there exists no contiguous guard $(g,[x,j+1])$ with $g\in\mathcal{E}[i-1,j+2]$. We therefore apply Lemma˜6.9 using the lookahead queries of the Sliding-Window data structure (3) to test in $O(\log n)$ time whether a guard $(g,[x,j+1])$ with $g\in\mathcal{E}[i-1,j+2]$ exists, and, if so, we increment $j$. This way, we find the pair $(i,j)$ that corresponds to $x,\textnormal{{next}}(x)$.
@@
We compute the shortest path from $x$ to $j+1$ in $O(\log n)$ time and obtain the supporting line $\ell$ of its first edge. A ray-shooting query in Sliding-Window data structure (3) identifies the last point of intersection $g^{*}$ between $\ell$ and $\mathcal{E}[i-1,j+1]$. By definition, the only ugly dominator that sees $x,\textnormal{{next}}(x)]$ has $g^{*}$ as its guard. We apply Lemma˜6.1 to compute the maximal chain $u_{\max},v_{\max}]$ that contains $j$ and is visible from $g^{*}$ in $O(\log n)$ time.
It follows that $\textnormal{{next}}(x)$ is realized either by $(g,[u,v])$ or by $(g^{*},[u_{\max},v_{\max}])$, and we can distinguish between these cases in constant time. We dequeue $x$ and proceed to the next element in the queue. Each $O(\log n)$ step is charged either to an increment of $j$ or to the removal of one element from the queue. Hence, the total running time is $O((n+|X|)\log n)$.
Theorem 7.3.
@@
Proof 7.4.
Let $X$ be the set that contains all vertices of $P$ and includes for all guards $(g,[u,v])\in\mathcal{D}^{\prime}$ the point $u$. Theorem˜6.7 implies $X$ can be computed in $O(n\log n)$ time. By Theorem˜4.17 and the proof of Theorem˜4.19, it follows that if we apply $\textnormal{{next}}(x)$ to each $x\in X$, recursively, $k$ times then for at least one $x\in X$ the result is a point $v>x+n$. In other words, for this $x\in X$ the recursive application of $\textnormal{{next}}(x)$ results in a solution of size $k$ that guards $\partial P$.
We first compute $k$ and such a corresponding $x\in X$, and afterwards find the corresponding optimal guarding solution. Via Lemma˜7.1, we can compute $\textnormal{{next}}(x)$ for every $x$ in $X$. We do this repeatedly, computing, and maintaining, $X_{i}=\{\textnormal{{next}}(x)|x\in X_{i-1}\}$, with $X_{0}=X$, until there is a point $x$ such that $\textnormal{{next}}^{k}(x)\geq x+n$. This takes $k$ rounds, and thus a total of $O(kn\log n)$ time, and returns the size $k$ of the optimal solution and some $x$. To also output the corresponding set of guards, we apply Lemma˜7.1 once more with an input $X=\{x\}$ where each time we find a guard $(g,[x,\textnormal{{next}}(x)])$, we quickly add $\textnormal{{next}}(x)$ to $X$ and the queue before we dequeue $x$. Since Lemma˜7.1 computes the explicit guard $(g,[x,\textnormal{{next}}(x)])$, the result are $k$ guards that together see all of $\partial P$.
For the remainder of this paper we will assume that we know that there is no solution of size $3$ or less. This allows us to impose stronger structure, which in turn enables us to give an $O(n\log n)$ algorithm for a polygon where only solutions of size at least $4$ exist. To this end, we show that Theorem˜7.3 implies an $O(n\log n)$-time test for this property:
Corollary 7.5.
@@
Proof 7.6.
We simply stop the algorithm described in Theorem˜7.3 after $K$ rounds.